8.) P 8500 is borrowed for 70 days at 18% per annum. How much will be due at the end of 70 days? Use exact simple interest.


1
Expert's answer
2022-03-09T16:25:02-0500

A = P0 + I = P0 + P0r = P0 (1 + r)

I is the interest;

A is the end amount: principal plus interest;

P0 is the principal (starting amount);

r is the interest rate (in decimal form)

r = 18% : 365 days × 70 days = 3.45% or 0.0345

A = 8500 (1 + 0.0345) = 8793